### v1

@@ 1 Introduction
where $\Omega\subset\mathcal{R}^{2}$ is a convex polygonal domain with a sufficiently smooth boundary $\partial\Omega$, the density of the fluid is denoted by $\rho=\rho(\mathbf{x},t)$, the velocity of the fluid is represented by $u=(u_{1}(\mathbf{x},t),u_{2}(\mathbf{x},t))^{\top}$, $\mu$ denotes the viscosity coefficient, $f=(f_{1}(\mathbf{x},t),f_{2}(\mathbf{x},t))^{\top}$ is a given function. Moreover, we give the following initial conditions and boundary conditions:
@@
The functions $\rho_{0}(\mathbf{x})$, $a(\mathbf{x},t)$, $u_{0}(\mathbf{x})$ and $g(\mathbf{x},t)$ are provided, the inflow boundary is defined as $\Gamma_{in}=\{\mathbf{x}\in\partial\Omega:g\cdot\vec{\nu}<0\}$, where $\vec{\nu}$ represents the outward normal vector, and the initial density $\rho_{0}(\mathbf{x})$ satisfy the following conditions [22]
@@
For the existence and uniqueness of the solutions of Navier-Stokes equations with variable density (1.1)-(1.3), the reader is referred to, e.g., [5, 9, 14, 29]. On the other hand, there have been lots of attentions in developing efficient numerical methods for (1.1)-(1.3), especially in the schemes preserving physical properties. In 1992, Bell et al. [2] first introduced the projection method for variable density issues, they employed the Crank-Nicolson method for temporal discretization, and utilized a standard difference method for spatial discretization. Subsequently, Almgren et al. [1] and Puckett et al. [33] investigated the conservative adaptive projection method and the higher-order projection method for tracking fluid interfaces, respectively. Unlike other traditional algorithms, this method reduces computational costs by solving the discrete pressure variable through the incorporation of a Poisson equation. In [20], a novel time-stepping method was introduced which had been verified by some numerical examples. Additionally, Li et al. in [19] proposed a second-order mixed stabilized finite element method for solving Navier-Stokes equations with variable density. Furthermore, faced with the same issue, Liu and Walkington [23] conducted an investigation into the discontinuous Galerkin (DG) method. They proved the convergence of the scheme but did not provide any convergence rates. In contrast, Pyo and Shen [34] studied two Gauge-Uzawa schemes and demonstrated that the first-order temporally discretized Gauge-Uzawa schemes possess unconditional stability. Moreover, Li and Wu [18] presented a filtered time-stepping technique [6], which could improve the time accuracy to second-order. Afterwards, Reuter et al.[35] introduced a novel algorithm of explicit temporal discretization for low-Mach Navier-Stokes equations with variable density, which achieved second-order accuracy in time. By constructing an implicit temporal scheme with the Taylor series and using a finite element with standard high-order Lagrange basis functions, Lundgren et al. [25] considered a fourth-order method for (1.1)-(1.3).
@@
When designing numerical schemes, one of interesting and challenging topics is to preserve the physical properties of the continuous model at the discrete scheme, which has attracted lots of attentions in the past decade. For the Navier-Stokes equations with constant density, by transforming into an equivalent form known as the EMAC formulation in [4], a mixed finite element method are proposed, which imposed the incompressible condition weakly and preserved physical properties such as momentum, energy, and enstrophy. This research was further extended to address long-term approximations in [28] and three-dimensional problems in [13]. Concurrently, a mimetic spectral element method was introduced in [30], that is capable of preserving mass, energy, enstrophy, and vorticity. Additionally, this concept was adapted to problems involving moving domains in [11]. Lately, by deriving the viscosity coefficients through a residual-based shock-capturing approach, Lundgren et al. [24] presented a novel symmetric and tensor-based viscosity method, which can ensure the conservation of angular momentum and the dissipation of kinetic energy. For the variable density incompressible flows, an entropy-stable scheme was explored in [27] by combining the discontinuous Galerkin method with an artificial compressible approximation. Recognizing the significance of density bounds in numerical simulations, a bound-preserving discontinuous Galerkin method was introduced in [17]. Furthermore, Desmons et al. [7] introduced a generalized high-order momentum preserving scheme, which was acknowledged to be easily implementable utilizing the finite volume method. To ensure the positivity preserving of the density, a square transformation $\rho=\sigma^{2}$ was introduced in [22, 21]. By introducing power-type and exponential-type scalar auxiliary variables to define the system’s energy and to balance the incompressible condition’s influence respectively, Zhang et al. [41] reformulated the Navier-Stokes equations with variable density into an equivalent form and subsequently developed a linear, decoupled, and fully discrete finite element scheme. This scheme preserves the mass, momentum, and modified energy conservation relations. Recently, by introducing a formulation with consistent nonlinear terms, the schemes with the numerical density invariant to global shifts was studied in [26]. And the authors in [16] investigate schemes which could preserve the lower bound of the numerical density and energy inequality under the gravitational force.
@@
But, due to the complex nonlinearities and coupling terms, it is challenging to derive error analysis for numerical methods solving the Navier-Stokes equations with variable density. Under the assumptions that the numerical density is bound and can achieves first order convergence, the author in [8] presented a first-order splitting scheme and deduced its error estimates. Recently, giving up the assumption on the numerical density, Cai et al. [3] derived the error estimate of the backward Euler method applied to the 2D Navier-Stokes equations with variable density, leveraging an error splitting technique and discrete maximal $L^{p}$-regularity. Drawing upon this research, Li and An in [22] presented a novel BDF2 finite element scheme, by utilizing the mini element space to approximate both the velocity and the pressure, and employing the quadratic conforming finite element space to approximate the density. Leveraging a post-processed technique, the authors in [15] demonstrated the convergence order of $O(\tau^{2}+h^{2})$ in $L^{2}$-norm for the numerical density $\rho_{h}^{n}$ and numerical velocity $u_{h}^{n}$. Lately, by rewriting the original system, Pan and Cai in [31] proposed a general BDF2 finite element method preserving the energy inequality and deduced its error analysis. But, there is no literature on error estimates for the fully discrete first-order scheme for the Navier-Stokes equations with variable density, which can preserve the positivity of the numerical density and the original energy identical-relation.
@@
In this paper, we will consider a mass conservation, positivity and energy identical-relation preserving scheme for the Navier-Stokes equations with variable density (1.1)-(1.3). To ensure the positivity of the numerical density, we utilize the square transformation considered in [22, 21] to transform the density sub-equation. Compared to other positivity preserving methods, the method considered here has two mainly advantages: form-invariant and irrelevance of the discrete scheme. Therefore, it is possible to directly adopt other schemes in the references for solving the density sub-equation. But, the mass conservation is lost when approximating this reformation form. To overcome this problem, then we use the recovery technique in [38] to preserve the discrete system’s mass. Moreover, to eliminate the numerical dissipation usually existent in the numerical scheme, we propose a new recovery method, which results that this scheme considered in this paper not only can inherit the mass conservation, positivity, original energy identical-relation from the continuous equations, but also achieve the following convergence order in the $L^{2}$-norm
@@ 2.1 Functional spaces
$||u||_{L^{p}(\Omega)}$ $=\left(\int_{\Omega}|u(\mathbf{x})|^{p}\,\mathrm{d}x\right)^{\frac{1}{p}},$
@@
Within this context, $W^{k,2}(\Omega)$ is also known as the Hilbert space and can be expressed as $H^{k}(\Omega)$. $||\cdot||_{L^{\infty}}$ represent the norm of space of $L^{\infty}(\Omega)$,
@@
On the other hand, let $\mathcal{T}_{h}=\{K\}$ be a uniformly regular triangulation partition of $\Omega$ with a mesh size $h(0<h<1)$. We also define the finite element space
$V_{h}=\{u_{h}\in C(\bar{\Omega})\cap V,~{}v_{h}|_{K}\in P_{2}(K)^{2},~{}\forall K\in\mathcal{T}_{h}\}\subset V,$
@@ 2.2 Some inequalities
We recall some useful inequalities in two dimension in this subsection. For any $v_{h}$ belongs to the finite element space defined above, there hold
@@ 2. Agmon’s inequality [10]:
$||v_{h}||_{L^{\infty}}\leq C||v_{h}||_{L^{2}}^{\frac{1}{2}}||\triangle v_{h}||_{L^{2}}^{\frac{1}{2}}.$ (2.4)
@@
Lemma 2.2.
[…]
$\left(\sum\limits_{K\in\mathcal{T}_{h}}||v-I_{h}v||_{W_{p}^{s}(K)}^{p}\right)^{\frac{1}{p}}\leq Ch^{m-s}|v|_{W_{p}^{m}(\Omega)}.$ (2.12)
@@ 2.3 Some essential properties
For the Navier-Stokes equations with variable density (1.1)-(1.3), there hold the following essential properties (see [41]):
@@ 3. Energy identical-relation:
To preserve the positivity, we adopt the square transformation [21]
@@
which guarantees that the density is non-negative regardless of the discrete scheme. Moreover, we can derive an equivalent formulation of the continuous equation below
@@
$\sigma(\sigma u)_{t}-\mu\triangle u+\rho(u\cdot\nabla)u+\frac{u}{2}\nabla\cdot(\rho u)+\nabla p$ $=f,$ $\mathrm{in}~{}\Omega\times(0,T],$ (2.15)
@@
$\sigma_{0}(\mathbf{x})=\sqrt{\rho_{0}(\mathbf{x})}\quad\mathrm{and}\quad\sqrt{\rho_{0}^{min}}\leq\sigma(t,\mathbf{x})\leq\sqrt{\rho_{0}^{max}},\quad~{}\mathrm{in}~{}\Omega,$ (2.17)
@@
The solution of (2.14)-(2.16) satisfies the following regularities [21, 22]:
@@
$u$ $\in C([0,T];H^{3}(\Omega))\cap C^{1}([0,T];H^{2}(\Omega)),~{}~{}~{}~{}p\in C([0,T];H^{2}(\Omega)).$
@@ 3 Property-preserving scheme
In this section, based on a new recovery technique, we will propose a fully discrete first order finite element scheme for solving the incompressible Navier-Stokes equations with variable density. Let $N\in\mathbb{N}^{+}$ and $\tau=T/N(0<\tau<1)$, thus $0=t_{0}<t_{1}<\dots<t_{k}<t_{k+1}\dots<t_{N}=T$. Define $D_{\tau}g^{n+1}:=\frac{g^{n+1}-g^{n}}{\tau}$, then the first order scheme considered in this paper is as follows: Given $(\sigma_{h}^{0},\rho_{h}^{0},u_{h}^{0},p_{h}^{0})=(\Pi_{h}\sigma^{0},\Pi_{h}\rho^{0},R_{h}u^{0},Q_{h}p^{0})$, find $(\sigma_{h}^{n+1},\tilde{u}_{h}^{n+1},p_{h}^{n+1},u_{h}^{n+1},\rho_{h}^{n+1})$ for $10\leq n\leq N-$ 1 through the following steps:
@@
$(D_{\tau}\sigma_{h}^{n+1},r_{h})+(\nabla\sigma_{h}^{n+1}\cdot u_{h}^{n},r_{h})+\frac{1}{2}(\sigma_{h}^{n+1}\nabla\cdot u_{h}^{n},r_{h})=0;$ (3.1)
@@
$(\sigma_{h}^{n+1}D_{\tau}(\sigma_{h}^{n+1}\tilde{u}_{h}^{n+1}),v_{h})+\mu(\nabla\tilde{u}_{h}^{n+1},\nabla v_{h})+(\rho_{h}^{n}(u_{h}^{n}\cdot\nabla)\tilde{u}_{h}^{n+1},v_{h})$ (3.2)
$+\frac{1}{2}(\tilde{u}_{h}^{n+1}\nabla\cdot(\rho_{h}^{n}u_{h}^{n}),v_{h})-(p_{h}^{n+1},\nabla\cdot v_{h})+(\nabla\cdot\tilde{u}_{h}^{n+1},q_{h})=(f^{n+1},v_{h});$
@@
In Steps 1-2, we get the approximation solutions $\sigma_{h}^{n+1},~{}\tilde{u}_{h}^{n+1}$ and $p_{h}^{n+1}$ by solving two linear system. But, the mass conservation and the original energy identical-relation is lost in Steps 1 and 2, respectively. To make these numerical solutions to satisfy the property of the continuous equations, we recover them in Steps 3-4, which are made up of several assignment operations and can be implemented efficiently. For the scheme (3.1)-(3.7), there hold the following Theorem.
@@
The scheme (3.1)-(3.7) inherits the following physical properties of the continuous equations (1.1)-(1.3):
1. Positivity: $\rho_{h}^{n+1}=\lambda_{h}^{n+1}\bar{\rho}_{h}^{n+1}>0$.
2. Mass conservation: $\int_{\Omega}\rho_{h}^{n+1}\,\mathrm{d}\mathbf{x}=\int_{\Omega}\rho_{h}^{n}\,\mathrm{d}\mathbf{x}=\int_{\Omega}\rho_{h}^{0}\,\mathrm{d}\mathbf{x}$.

### v2

@@ 1 Introduction
where $\Omega\subset\mathcal{R}^{2}$ is a convex polygonal domain with a sufficiently smooth boundary $\partial\Omega$, $\rho=\rho(\mathbf{x},t)$ represents the density of the fluid, $u=u(\mathbf{x},t)=(u_{1}(\mathbf{x},t),u_{2}(\mathbf{x},t))^{\top}$ represents the velocity of the fluid, $\mu$ denotes the viscosity coefficient, $f=(f_{1}(\mathbf{x},t),f_{2}(\mathbf{x},t))^{\top}$ is a given body force. Moreover, we give the following initial conditions and boundary conditions:
@@
$\rho_{0}(\mathbf{x})$, $a(\mathbf{x},t)$, $u_{0}(\mathbf{x})$ and $g(\mathbf{x},t)$ are given functions, $\Gamma_{in}=\{\mathbf{x}\in\partial\Omega:g\cdot\vec{\nu}<0\}$ is the inflow boundary with $\vec{\nu}$ being the outward normal vector, and the initial density $\rho_{0}(\mathbf{x})$ satisfy the following conditions [24]
@@
For the existence and uniqueness of the solutions of Navier-Stokes equations with variable density (1.1)-(1.3), the reader is referred to, e.g., [5, 9, 15, 31]. On the other hand, there have been lots of attentions in developing efficient numerical methods for (1.1)-(1.3), especially in the schemes preserving physical properties. In 1992, Bell et al. [2] first introduced the projection method for variable density issues, they employed the Crank-Nicolson method for temporal discretization, and utilized a standard difference method for spatial discretization. Subsequently, Almgren et al. [1] and Puckett et al. [35] investigated the conservative adaptive projection method and the higher-order projection method for tracking fluid interfaces, respectively. Unlike other traditional algorithms, this method reduces computational costs by solving the discrete pressure variable through the incorporation of a Poisson equation. In [22], a novel time-stepping method was introduced which had been verified by some numerical examples. Additionally, Li et al. in [21] proposed a second-order mixed stabilized finite element method for solving Navier-Stokes equations with variable density. Furthermore, Liu and Walkington [25] conducted an investigation into the discontinuous Galerkin (DG) method for solving Navier-Stokes equations with variable density. They proved the convergence of the scheme but did not provide any convergence rates. In contrast, Pyo and Shen [36] studied two Gauge-Uzawa schemes and demonstrated that the first-order temporally discretized Gauge-Uzawa schemes possess unconditional stability. Moreover, Li et al. [19] presented a filtered time-stepping technique [6], which could improve the time accuracy to second-order. Afterwards, Reuter et al. [38] introduced a novel algorithm of explicit temporal discretization for low-Mach Navier-Stokes equations with variable density, which achieved second-order accuracy in time. By constructing an implicit temporal scheme with the Taylor series and using a finite element with standard high-order Lagrange basis functions, Lundgren et al. [27] considered a fourth-order method for (1.1)-(1.3).
@@
When designing numerical schemes, one of interesting and challenging topics is to preserve the physical properties of the continuous model in the discrete scheme, which has attracted lots of attentions in the past decade. For the Navier-Stokes equations with constant density, by transforming into an equivalent form known as the energy, momentum and angular momentum conserving (EMAC) formulation in [4], a mixed finite element method are proposed, which imposed the incompressible condition weakly and preserved physical properties such as momentum, energy, and enstrophy. This research was further extended to address long-term approximations in [30] and three-dimensional problems in [14]. Concurrently, a mimetic spectral element method was introduced in [32], that is capable of preserving mass, energy, enstrophy, and vorticity. Additionally, this concept was adapted to problems involving moving domains in [11]. Lately, by deriving the viscosity coefficients through a residual-based shock-capturing approach, Lundgren et al. [26] presented a novel symmetric and tensor-based viscosity method, which can ensure the conservation of angular momentum and the dissipation of kinetic energy. For the variable density incompressible flows, an entropy-stable scheme was explored in [29] by combining the discontinuous Galerkin method with an artificial compressible approximation. Recognizing the significance of density bounds in numerical simulations, a bound-preserving discontinuous Galerkin method was introduced in [18]. Furthermore, Desmons et al. [7] introduced a generalized high-order momentum preserving scheme, which was claimed to be easy for implementation with the finite volume method. To ensure the positivity preserving of the density, a square transformation $\rho=\sigma^{2}$ was introduced in [23, 36, 42]. By introducing power-type and exponential-type scalar auxiliary variables to define the system’s energy and to balance the incompressible condition’s influence respectively, Zhang et al. [44] reformulated the Navier-Stokes equations with variable density into an equivalent form and subsequently developed a linear, decoupled, and fully discrete finite element scheme. This scheme preserves the mass, momentum, and modified energy conservation relations. Recently, by introducing a formulation with consistent nonlinear terms, the schemes with the numerical density invariant to global shifts was studied in [28]. And the authors in [17] investigate schemes which could preserve the lower bound of the numerical density and energy inequality under the gravitational force.
@@
But, due to the complex nonlinearities and coupling terms, it is challenging to derive error analysis for numerical methods solving the Navier-Stokes equations with variable density. Under the assumptions that the numerical density is bound and can achieves first order convergence, the author in [8] presented a first-order splitting scheme and deduced its error estimates. Recently, giving up the assumption on the numerical density, Cai et al. [3] derived the error estimate of the backward Euler method applied to the 2D Navier-Stokes equations with variable density, leveraging an error splitting technique and discrete maximal $L^{p}$-regularity. Drawing upon this research, Li and An in [24] presented a novel BDF2 finite element scheme, by utilizing the Mini element space to approximate both the velocity and the pressure, and employing the quadratic conforming finite element space to approximate the density. Leveraging a post-processed technique, the authors in [16] demonstrated the convergence order of $O(\tau^{2}+h^{2})$ in $L^{2}$-norm for the numerical density $\rho_{h}^{n}$ and numerical velocity $u_{h}^{n}$. Lately, by rewriting the original system, Pan and Cai in [33] proposed a general BDF2 finite element method preserving the energy inequality and deduced its error analysis. But, there is no literature on error estimates for the fully discrete first-order scheme for solving Navier-Stokes equations with variable density, which can preserve the mass conservation, the positivity of the numerical density and the original energy identical-relation of the system.
@@
In this paper, we will consider a mass conservation, positivity and energy identical-relation preserving scheme for the Navier-Stokes equations with variable density (1.1)-(1.3). To ensure the positivity of the numerical density, we utilize the square transformation considered in [23, 42] to transform the density sub-equation. Compared to other positivity preserving methods, the method considered here has two mainly advantages: form-invariant and irrelevance of the discrete scheme. Therefore, it is possible to directly adopt other schemes in the references for solving the density sub-equation. But, the mass conservation is lost when approximating this reformation form. To overcome this problem, then we use the recovery technique in [12, 41] to preserve the discrete system’s mass. In addition, through constructing a new recovery method, we eliminate successfully the numerical energy dissipation usually existent in the numerical scheme. Moreover, we prove that the scheme considered in this paper not only can inherit the mass conservation, positivity, original energy identical-relation from the continuous equations, but also achieve the following convergence order in the $L^{2}$-norm
@@ 2.1 Functional spaces
$||u||_{L^{p}(\Omega)}$ $=\left(\int_{\Omega}|u(\mathbf{x})|^{p}\,\mathrm{d}\mathbf{x}\right)^{\frac{1}{p}},$
@@
Within this context, $W^{k,2}(\Omega)$ is also known as the Hilbert space and can be expressed as $H^{k}(\Omega)$. $||\cdot||_{L^{\infty}}$ represents the norm of the space $L^{\infty}(\Omega)$ which is defined as
@@
On the other hand, let $\mathcal{T}_{h}=\{K\}$ be a uniformly regular triangulation partition of $\Omega$ with a mesh size $h(0<h<1)$. We also define the finite element spaces
$V_{h}=\{u_{h}\in C(\bar{\Omega})^{2}\cap V,~{}v_{h}|_{K}\in P_{2}(K)^{2},~{}\forall K\in\mathcal{T}_{h}\}\subset V,$
@@ 2.2 Some inequalities
We recall some useful inequalities in two dimension in this subsection. For any $v_{h}$ belongs to the finite element spaces defined above, there hold
@@ 2. Agmon’s inequality [10]:
$||v_{h}||_{L^{\infty}}\leq C||v_{h}||_{L^{2}}^{\frac{1}{2}}||\Delta v_{h}||_{L^{2}}^{\frac{1}{2}}.$ (2.4)
@@ 2.3 Some essential properties
For the Navier-Stokes equations with variable density (1.1)-(1.3), there hold the following essential properties (see, i.e., [23, 24, 36, 44]):
@@ 3. Energy identical-relation:
To preserve the positivity, we adopt the square transformation [23, 36, 42]
@@
which guarantees that the density is non-negative regardless of the discrete scheme. Moreover, to derive the energy relation of the considered scheme, we adopt an equivalent formulation of the momentum equation (1.2) (see, i.e., [36, 42]), which combining with (2.12) and (1.3) yields
@@
$\sigma(\sigma u)_{t}-\mu\Delta u+\rho(u\cdot\nabla)u+\frac{u}{2}\nabla\cdot(\rho u)+\nabla p$ $=f,$ $\mathrm{in}~{}\Omega\times(0,T],$ (2.14)
@@
$\sigma_{0}(\mathbf{x})=\sqrt{\rho_{0}(\mathbf{x})}\quad\mathrm{and}\quad 0<\sqrt{\rho_{0}^{min}}\leq\sigma(t,\mathbf{x})\leq\sqrt{\rho_{0}^{max}},\quad~{}\mathrm{in}~{}\Omega,$ (2.16)
@@
The solutions of (2.13)-(2.15) satisfy the following regularities [23, 24]:
@@
$u$ $\in C([0,T];H^{3}(\Omega)^{2})\cap C^{1}([0,T];H^{2}(\Omega)^{2}),~{}~{}~{}~{}p\in C([0,T];H^{2}(\Omega)).$
@@ 3 Property-preserving scheme
In this section, we will propose a property-preserving fully discrete first order finite element method for solving the incompressible Navier-Stokes equations (2.13)-(2.15) with variable density. Although the positivity of the density is preserved by using the square transformation (2.12), the mass conservation will be lost when approximating this reformation form. Adopting the recovery technique in [12, 41], we recovery the discrete system’s mass. In addition, via constructing a new recovery method, we also eliminate the numerical energy dissipation which is usually existent in the classical scheme, which ensures the energy identical-relation of the proposed scheme.
Let $N\in\mathbb{N}^{+}$ and $\tau=T/N(0<\tau<1)$, thus $0=t_{0}<t_{1}<\dots<t_{k}<t_{k+1}\dots<t_{N}=T$. Define $D_{\tau}g^{n+1}:=\frac{g^{n+1}-g^{n}}{\tau}$, then the first order scheme for the equations (2.13)-(2.15) considered in this paper is as follows: Given $(\sigma_{h}^{0},\rho_{h}^{0},u_{h}^{0})=(\Pi_{h}\sigma^{0},\Pi_{h}\rho^{0},R_{h}u^{0})$, find $(\sigma_{h}^{n+1},\tilde{u}_{h}^{n+1},p_{h}^{n+1},u_{h}^{n+1},\rho_{h}^{n+1})$ for $10\leq n\leq N-$ 1 through the following steps:
@@
$(D_{\tau}\sigma_{h}^{n+1},r_{h})+(\nabla\sigma_{h}^{n+1}\cdot u_{h}^{n},r_{h})+\frac{1}{2}(\sigma_{h}^{n+1}\nabla\cdot u_{h}^{n},r_{h})=0,\quad\forall r_{h}\in W_{h};$ (3.1)
@@
$(\sigma_{h}^{n+1}D_{\tau}(\sigma_{h}^{n+1}\tilde{u}_{h}^{n+1}),v_{h})+\mu(\nabla\tilde{u}_{h}^{n+1},\nabla v_{h})+(\rho_{h}^{n}(u_{h}^{n}\cdot\nabla)\tilde{u}_{h}^{n+1},v_{h})+\frac{1}{2}(\tilde{u}_{h}^{n+1}\nabla\cdot(\rho_{h}^{n}u_{h}^{n}),v_{h})$ (3.2)
$-(p_{h}^{n+1},\nabla\cdot v_{h})+(\nabla\cdot\tilde{u}_{h}^{n+1},q_{h})=(f^{n+1},v_{h}),\quad\forall(v_{h},q_{h})\in(V_{h},M_{h});$
@@
In Steps 1-2, we get the approximation solutions $\sigma_{h}^{n+1},~{}\tilde{u}_{h}^{n+1}$ and $p_{h}^{n+1}$ by solving two linear system. But, the mass conservation and the original energy identical-relation is lost in Steps 1 and 2, respectively. To make the scheme to satisfy the property of the continuous equations, we recover them in Steps 3-4, which are made up of several assignment operations and can be implemented efficiently. For the scheme (3.1)-(3.7), there holds the following Theorem.
@@
The scheme (3.1)-(3.7) inherits the following physical properties of the continuous equations (1.1)-(1.3) for $10\leq n\leq N-$ 1:
1. Positivity: $\rho_{h}^{n+1}>0$.
2. Mass conservation: $\int_{\Omega}\rho_{h}^{n+1}\,\mathrm{d}\mathbf{x}=\int_{\Omega}\rho_{h}^{0}\,\mathrm{d}\mathbf{x}$.